OpenStudy (anonymous):

Find the hypotenuse of a right triangle with legs of 3 inches and 4 inches...... I got 7 but i am wrong please help

OpenStudy (anonymous):

do 3sq+4sq=csq (sorry i couldnt find the carat

OpenStudy (anonymous):

which becomes 9+16=csq

OpenStudy (anonymous):

25=csq

OpenStudy (anonymous):

then you find the square root of both sides

OpenStudy (amistre64):

shift+6 = ^

OpenStudy (anonymous):

lol thx

OpenStudy (anonymous):

so is it 5?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

wooooooooo

OpenStudy (anonymous):

thxn guys

OpenStudy (amistre64):

\[3+4=7\] but that aint pythag \[3^2+4^2=c^2\]

OpenStudy (anonymous):

Find the hypotenuse of a right triangle with legs of 9 inches and 12 inches= 12?